A tungsten filament lamp has a resistance of 11.8 ohms at 20 degrees Celsius. Determine its resistance at 2500 degrees Celsius


1
Expert's answer
2021-02-09T10:41:53-0500

Calculate the value of the difference in the temperature.

"ΔT = T_2 -T_1 \\

= 2500 -20 \\

= 2480 \;ºC"

Consider the expression for the new value of the temperature due to increase in the temperature.

"R_2 =R_1(1 + αΔT) \\

= 11.8 (1 + 0.0045 \times 2480) \\

= 11.8 (1 +11.16) \\

= 11.8 \times 12.16 \\

= 143.5 \;Ω"

Answer: 143.5 Ω